OpenStudy (anonymous):

Simplify The Following, Show work. {sqrt192} {Sqrt147} {sqrt98} {sqrt44} {sqrt116} {sqrt171} {sqrt63} ((TY))

myininaya (myininaya):

I think it might be easier to write prime factorizations for the numbers in the square roots first

OpenStudy (anonymous):

me too in fact, i think that is the only way to do it annoying, but what can you do when your math teacher hates you?

OpenStudy (anonymous):

for example \[192=2^6\times 3\]

OpenStudy (anonymous):

\[147=3\times 7^2\]

OpenStudy (anonymous):

etc want to know how i get them so fast?